# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v1

@@ III.5 Polycatenanes: Steady Uniaxial Extensional Flow
As shown in Fig. 6b, increasing the ring size while keeping the number of rings constant leads to more variability of $\langle X_{bb}\rangle/L$ because the beads in the same ring experience less repulsions until they become fully extended. The chains with larger rings have more flexibility Rauscher et al. (2020) and can collapse and extend more readily. As a result, in weak flows, chains with smaller rings exhibit higher $\langle X_{bb}\rangle/L$, whereas in strong flows, chains with larger rings show greater extension.
@@
According to Fig. 6c, at low Wi, polycatenanes have higher $(N_{1})$ than linear polymers with the same molecular weight because the interlocks prevent chain collapse and increase stress buildup. At medium to high Wi, linear polymers are more extended and have higher $(N_{1})$. Furthermore, $(N_{1})$ increases with increasing number of rings and ring size primarily due to higher $M_{w}$. Furthermore, $(N_{1})$ increases with increasing number of rings primarily due to higher $M_{w}$. Increasing ring size has a similar effect, as evident from Fig. 6d.
@@
Extensional viscosity ($\eta_{e}$) is significantly lower in polycatenanes compared to linear polymers with the same molecular weight, as shown in Fig. 6e. For instance, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) has 22 times lower ($\eta_{e}$) than a linear polymer with 80 beads at $\mathrm{Wi}=10$. Furthermore, adding more rings to a polycatenane increases the extensional viscosity due to increased molecular weight of the polymers. The same effect is achieved by making the rings larger, as seen in Fig. 6f. Comparing a polycatenane with 16 beads per ring and 10 rings ($M_{w}=160$) to a polycatenane with 8 beads per ring and 21 rings ($M_{w}=168$) allows us to decouple the effects of molecular weight from the impact of chain topology on viscosity. Based on Figs. 6e and 6f, at low Wi, the chain with more small rings has higher viscosity. This reflects the more extended conformation of the smaller rings, leading to higher stress and resistance to extension. At high Wi, the chain with fewer large rings show significantly higher extensional viscosity, up to 270% higher at the highest Wi. Under strong extensional flow, the rings in the chain with fewer large rings are more effectively stretched, leading to an increased resistance to further deformation. The extended conformation under high strain produces a larger stress response, which is reflected in the higher measured extensional viscosity. In contrast, the greater number of flexible mechanical interlocks in the chain with more small rings facilitates stress relaxation when subjected to high strain rates. This relaxation mechanism ultimately reduces $\eta_{e}$ relative to the chain with fewer large rings.
III.6 Polycatenanes: Steady Shear Flow
@@
As shown in Fig. 7a, at low Wi, $\langle X_{bb}\rangle/L$ remains nearly constant in polycatenanes as the flow is not strong enough to create significant deformation in the chains. For $1\leq\mathrm{Wi}\leq 100$, there is a regime with increasing extension which corresponds to chains being extended in flow direction. In chains with 21 or more rings, there is a decreasing regime in the $100\leq\mathrm{Wi}$ range. This corresponds to expansion in cross-flow direction. This behavior could not be verified in chains with 5 and 10 rings because the chains could not withstand the strength of the flow and dethreaded. At low Wi, $\langle X_{bb}\rangle/L$ is higher in polycatenanes than linear polymers with the same molecular weigth due to entangled ring topologies leading to strong repulsive interactions, keeping the chains more extended. At higher Wi, linear polymers become more extended in flow direction than the polycatenanes. This is because the linear polymers are free to change conformation from coiled to stretched states in the flow, while the rings in polycatenanes have steric constraints, preventing them from significant deformation. In addition, mechanical interlocks limit the extension of the chains due to inter-ring repulsion. Chains with fewer rings have limited variation of $\langle X_{bb}\rangle/L$ throughout the Wi range, while chains with higher number of rings can be deformed more readily in the flow, leading to larger variation of fractional extension in flow direction. In addition, higher number of rings leads to lower fractional extension as the chains are more flexible and can relax to a coiled state. Based on 7b, increasing the ring size has the same effect on $\langle X_{bb}\rangle/L$ at low to medium Wi. At high Wi, the effect becomes negligible.
Figs. B.3a and B.3b show that $\langle Y_{bb}\rangle/L$ and $\langle Z_{bb}\rangle/L$ are higher in polycatenanes than linear polymers because the rings add extra span in those directions, while increasing ring size or adding more rings decrease the extensions primarily due to the normalization factor effects. $\langle Y_{bb}\rangle/L$ and $\langle Z_{bb}\rangle/L$ show a decreasing regime, while $\langle Z_{bb}\rangle/L$ decreases only up to $\mathrm{Wi}=100$ after which an increasing regime is evident corresponding to decreasing $\langle X_{bb}\rangle/L$. According to Fig. B.3c, at low Wi polymers remain unaligned. At moderate Wi, the orientation angle $\theta$ first rises then falls, with polycatenanes of 8 beads per ring exhibiting a secondary increase previously noted in other systems Dutta and Sing (2024). A terminal decrease, reflecting alignment with very strong flow, could not be verified here because chains broke at high Wi. Having more or smaller rings shifts the primary peak to higher Wi and brings the secondary rise to lower Wi. Having more or larger rings make the orientation angle approach that of a linear polymer.
Polycatenanes show significantly lower shear viscosity than linear polymers, as shown in Fig. 7c. For example at $\mathrm{Wi}=10$, the viscosity of a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) is 35% of the linear polymer with 80 beads. This is due to the larger polymer extent in gradient direction which creates larger shear forces and larger deformation. In addition, viscosity varies less in the polycatenanes across the simulated Wi range compared with linear polymers. The power-law exponent for polycatenanes is 50% less than the linear polymers at the same molecular weight, suggesting weaker shear-thinning behavior. This can be understood by the fact that in linear polymers, higher shear rates cause the polymers to uncoil, extend, and align with the flow, leading to a significant decrease in viscosity. In contrast, polycatenanes do not deform, coil, or uncoil as readily under shear due to the repulsive interactions inherent in their interlocked structure, resulting in a flatter viscosity profile and less sensitivity to changes in shear rate. Increasing the number of rings in polycatenanes leads to a higher viscosity, primarily because of the increase in molecular weight. The slope of the viscosity curve increases as well. This is due to enhanced flexibility, leading to more prominent coiling and uncoiling, which leads to stronger shear-thinning properties. According to Fig. 7d, increasing the ring size has a similar effect, leading to higher shear viscosity and enhanced shear-thinning behavior.
@@
At very low Wi number, polycatenanes have decreased tumbling frequency compared to linear polymers, as shown in Fig. 7e. This is due to higher polymer extent in flow direction, creating greater moment of inertia, making it more difficult for a weak flow to rotate the chain. However, because of the interlocked rings, polycatenanes are thicker in gradient direction, making them more responsive to the flow strength. As a result, a slight increase in Wi causes more tumbling in polycatenanes compared with linear polymers, where tumbling frequency remains fairly constant in low Wi regime. At moderate to high Wi number, polycatenanes tumble more efficiently than linear polymers. For instance, at $\mathrm{Wi}=200$, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) tumbles 3 times more frequently than the linear polymer with 80 beads. Additionally, more rings in a polycatenane decreases tumbling frequency at moderate to high Wi number. This can be explained by the resulting increase in aspect ratio of the polymer, making the polymer thickness less effective, causing the chains to behave more like linear polymers. Increasing the ring size has a similar effect on tumbling frequency, as shown in Fig. 7f. This can be understood by the increased degrees of freedom at the mechanical interlocks, which allows the rings to rotate freely against each other and dissipate the shear stress, leading to lower tumbling frequency of the polymer.
@@ IV Discussion
Figure 8: Comparison of linear and MIP polymers of similar molecular weight in extensional (a, b) and shear (c, d) flow. Fractional extension in the flow direction $(\langle X_{bb}\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black) has 80 backbone beads ($M_{w}=80$). The polyrotaxane (blue) has 40 backbone beads including 2 capping beads and 6 rings ($M_{w}=88$). The daisy chain (green) has 4 linear segments of 8 beads each interlocked with 6 8 bead rings ($M_{w}=74$) and includes 2 capping beads. The polycatenane (red) has 10 interconnected rings of 8 beads each $M_{w}=80$.
@@
The presence of mechanical bonds gives rise to notable quantitative differences compared with linear polymers. Unlike linear polymers that coil under weak flow, in polyrotaxanes, ring-backbone repulsion causes more extended equilibrium configurations and earlier chain extension in flow. Polyrotaxanes show lower normal stress difference and weaker elasticity due to stress relief from sliding rings. Extensional viscosity is higher in polyrotaxanes especially in shorter chains with more rings due to increased stiffness and resistance to elongation. Under shear flow, polyrotaxanes extend more in all directions and have higher viscosity, with reduced shear thinning, as the rings limit mobility, requiring more stress for deformation. In addition, tumbling frequency is higher and it increases with number of rings due to enhanced hydrodynamic cross-section from threaded rings.
@@
Daisy chains extend less than linear polymers and have lower normal stress difference, extensional and shear viscosity. They tumble more frequently at high Wi, due to higher gradient-direction extent and aspect ratio effects with more or shorter segments pronouncing these effects, causing greater deviation from linear polymers.
@@
Polycatenanes are less extended and have lower first normal stress difference than linear chains except at very low Wi. They lack the smooth transition from coiled to stretched states typical of linear polymers. Polycatenanes show significantly lower extensional and shear viscosities, as well as weaker shear thinning. More or larger rings in polycatenanes increases viscosity, primarily due to increased molecular weight. Polycatenanes tumble more frequently, with fewer or smaller rings enhancing this effect.
@@
Overall, the interlocked topology of each MIP leads to a distinct set of rheological properties, as illustrated in Fig. 8. Here we give an example of each polymer type, all with a similar molecular weight. We compare polymer types with similar maximum contour lengths ($L$) in Fig. B.5. For a given molecular weight, we see that in both extensional (Fig. 8a) and shear flow (Fig. 8c), polyrotaxanes exhibit the greatest relative stretch in the flow direction because of their linear backbone and large cross-flow extent due to their rings. The daisy chain and the polycatenane are relatively unextended because of the lower extensibility of rings and their resistance to overlap. This leads to fairly consistent stretch of the daisy chain and the polycatenane across a wide range of Wi, varying only between 0.10 and 0.37 in extension and 0.09 and 0.18 in shear, much lower variability than other two polymer types. For a given molecular weight, the polymers also exhibit variation in their viscosity. In extensional flow, the linear polymer has the greatest extensional viscosity for all but the lowest Wi range, as seen in Fig. 8b. The mechanical bonds in all MIP types lead to lower extensional viscosity, with the daisy chain behaving similar to the linear polymer and the polyrotaxane and polycatenane showing weaker scaling of their extensional viscosity than the linear polymer and daisy chain. In shear flow (Fig. 8d), the linear polymer has higher shear viscosity than the MIPs except the polyrotaxane, which overtakes the linear polymer at a cross-over around $\mathrm{Wi}\approx 50$. This demonstrates the reduced shear thinning of the polyrotaxane. Indeed all of the MIPs’ shear viscosity scales less strongly with Wi than the linear polymer, demonstrating that the mechanical bonds suppress shear-thinning behavior. Similar to extensional viscosity, the shear viscosity of the polycatenane is the lowest of all polymer types.
@@
In line with the varied rheological responses that we have reported, MIPs have already shown their potential in a range of applications. Slide-ring gels, a category of polyrotaxanes in which mechanically interlocked rings function as movable cross-links, could potentially be used as protective coatings in automobiles and smartphones due to their scratch-resistance and self-healing properties Noda et al. (2014). Olympic gels (materials formed entirely by interlocked rings) Raphaël et al. (1997), which have been synthesized recently using over 16,000 DNA rings, exhibit unique non-linear elasticity and swelling behavior Speed et al. (2024). In addition, previous simulations of rotaxanes, slide-ring gels and tendomer gels in equilibrium and mechanical strain have demonstrated exceptional elasticity and yield stresses surpassing conventional cross-linked polymer gels Li et al. (2021); Yasuda et al. (2019); Chen et al. (2023); Tanahashi and Koga (2021); Uehara et al. (2022); Yasuda et al. (2020); Müller et al. (2019, 2021, 2022). Furthermore, the unusual physical characteristics of polysiloxanes along with the depolymerization behavior and elastic, rubber-like properties of polyphosphonitrilic chloride have been attributed to interlocked cyclic polymer structures Frisch et al. (1953); Patat and Derst (1959). These examples demonstrate the unique properties imparted by mechanical bonds. Our results hint at additional potential uses of MIPs, as well. High molecular weight polymers are known to reduce drag in pipe flow Larson (2003); Perkins et al. (1997); Dimitropoulos et al. (1998); Paterson and Abernathy (1970), at least partially due to the suppression of vortex formation caused by extended polymer configurations. The lower shear viscosity for MIPs compared to typical linear polymers and the greater extension of polyrotaxanes in particular, then, suggest that MIPs could be developed into drag reducing agents. Similarly, extended polymers in liquid petroleum-based fuels favor the formation of larger droplets that are less prone to ignition during accidental fuel releases Chao et al. (1984); Wei et al. (2015); Lhota et al. (2024). The greater extension of polyrotaxanes also suggests their potential development as mist-suppressing fuel additives to improve liquid fuel safety.
@@
As a first step toward characterizing the flow behavior of mechanically interlocked polymers, this work provides valuable insights into their dynamics in dilute solutions under weak to moderate flow conditions. However, strong flow regimes remain unexplored as our current model fails to maintain polymer integrity at high Weissenberg numbers. Under such conditions, large hydrodynamic forces can cause the rings in all three MIP architectures to slip through the polymer backbone, or through each other in the case of polycatenanes. This degradation is consistent with experiments that demonstrate that rotaxanes undergo unstoppering under high mechanical force Zhang and De Bo (2019). Experimental observations also suggest that mechanical bonds in catenanes are comparable in strength to covalent bonds Lee et al. (2016). To fully capture degradation mechanisms in strong flow, future work will move beyond the FENE bond model and incorporate dissociable bonds, revealing combined chemical and mechanical modes of polymer failure. This future work will lead to a better understanding of the durability of MIPs in flow for potential engineering applications.

### v2

@@ III.5 Polycatenanes: Extensional Flow
Fig. 6a shows that the variation of fractional extension of polycatenanes in the flow direction is less than linear polymers, and it is lower in polycatenanes that have fewer rings. For instance, a polycatenane with 8 beads per ring and 5 rings extends only 2.4 times beyond equilibrium, while a polycatenane chain with the same ring size and 65 rings stretches 4.2 times. More rings give the chain more degrees of freedom, leading to greater relative fluctuations in size and shape. In the limiting case, if we continue increasing the number of rings indefinitely, the chain approaches a linear polymer Brereton (2001). At high Wi, the difference becomes smaller as both polycatenanes and linear polymers approach fully extended configurations. As shown in Fig. 6b, increasing the ring size while keeping the number of rings constant also leads to more variability in extension due to the higher conformational degrees of freedom from the larger rings. The chains with larger rings have more flexibility Rauscher et al. (2020) and can collapse and extend more readily. As a result, in weak flows, chains with smaller rings exhibit higher $\langle X_{bb}\rangle/L$, whereas in strong flows, chains with larger rings show greater extension.
@@
According to Fig. 6c, at low Wi, polycatenanes have higher $N_{1}$ than linear polymers with the same molecular weight because the interlocks prevent chain collapse and increase stress. At medium to high Wi, linear polymers have higher $N_{1}$. Furthermore, $N_{1}$ increases with increasing number of rings because the polymer becomes subjected to larger forces from the extensional flow since there are more segments farther away from the stagnation point. Increasing ring size has a similar effect, as evident from Fig. 6d.
@@
Extensional viscosity ($\eta_{e}$) is significantly lower in polycatenanes compared to linear polymers with the same molecular weight, as shown in Fig. 6e. For instance, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) has 22 times lower ($\eta_{e}$) than a linear polymer with 80 beads at $\mathrm{Wi}=10$. Furthermore, adding more rings to a polycatenane increases the extensional viscosity due to increased molecular weight. The same effect is achieved by making the rings larger, as seen in Fig. 6f. Comparing a polycatenane with 16 beads per ring and 10 rings ($M_{w}=160$) to a polycatenane with 8 beads per ring and 21 rings ($M_{w}=168$) allows us to decouple the effects of molecular weight from the impact of chain topology on viscosity. Based on Figs. 6e and 6f, at low Wi, the chain with more small rings has higher viscosity. This reflects the more extended conformation of the smaller rings, leading to higher stress and resistance to extension. At high Wi, the chain with fewer large rings show significantly higher extensional viscosity, up to 270% higher at the highest Wi. Under strong extensional flow, the rings in the chain with fewer large rings are more effectively stretched, leading to an increased resistance to further deformation. The extended conformation under high strain produces a larger stress response, which is reflected in the higher measured extensional viscosity. In contrast, the greater number of flexible mechanical interlocks in the chain with more small rings facilitates lower stress response when subjected to high strain rates which ultimately reduces $\eta_{e}$ relative to the chain with fewer large rings.
III.6 Polycatenanes: Shear Flow
@@
As shown in Fig. 7a, at low Wi, polycatenanes exhibit a range of extensional behaviors based on the number of rings in their structure. Chains with fewer rings have limited variation of $\langle X_{bb}\rangle/L$ throughout the Wi range, while chains with higher number of rings can be deformed more readily in the flow, leading to larger variation of fractional extension in flow direction. For example, a polycatenane of 5 rings stretches between 0.32 and 0.43 of its maximum, compared to a 65 ring polycatenane, which extends between 0.17 and 0.35. Shorter, less flexible polycatenanes begin to stretch in the flow direction at lower flow strength than longer polycatenanes. For example, a 5-ring polycatenane shows increased stretch at $\mathrm{Wi}=0.1$, but is delayed to $\mathrm{Wi}=0.2$ for a 65-ring polycatenane. For $10^{0}<\mathrm{Wi}<\ 10^{2}$, there is a regime with increasing extension, which corresponds to chains being extended in flow direction with corresponding thinning in the gradient and cross-flow directions, shown in SI Figs. B.3a and B.3b, respectively. In chains with 21 or more rings, there is a decreasing regime in the $100\leq\mathrm{Wi}$ range. This corresponds to expansion in cross-flow direction, shown in SI Fig. B3.b. This behavior could not be verified in chains with 5 and 10 rings because of failure of the mechanical bonds due to bond crossing, a subject of future work. At low to moderate Wi, $\langle X_{bb}\rangle/L$ is higher in polycatenanes than linear polymers with the same molecular weight due to entangled ring topologies leading to strong repulsive interactions, keeping the chains more extended. or example, at $\mathrm{Wi}=1$, a 10-ring polycatenane has a fractional extension of 0.32 compared to a linear polymer of the same molecular weight ($M_{w}=80$) with an extension of 0.13. At high Wi, linear polymers are more extended, due to the more pronounced decreasing regime in polycatenanes. Based on 8a, increasing the ring size has the same effect on extension at low to medium Wi as number of rings. The greater flexibility of larger rings leads to more coiling behavior and less extended configurations.
Polycatenanes show significantly lower shear viscosity than linear polymers of similar molecular weight, as shown in Fig. 7b. For example at $\mathrm{Wi}=10$, the viscosity of a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) is 35% of the linear polymer with 80 beads. This decreased viscosity comes from the decreased size. Compared with linear polymers of the same molecular weight, polycatenanes have smaller size Lei et al. (2021); Pakula and Jeszka (1999), also demonstrated by equilibrium $R_{g}$ data in SI Table A.4. For example, a polycatenane with $M_{w}=80$ has $R_{g}=7.36$, about 57% of linear polymer with the same molecular weight, $R_{g}=12.85$. While their viscosity is lower, polycatenanes also have weakened shear-thinning. The viscosity power-law exponent for polycatenanes is 50% less than the linear polymers at the same molecular weight, shown in SI Table E.1. Consistent with both polyrotaxanes and daisy chains, polycatenanes also do not align as well with the flow as linear polymers, showcased by their larger orientation angles at high Wi in SI Fig. B.3c. Flow alignment is better for longer polycatenanes, leading to more shear thinning for these larger molecules. From Fig. 8b, increasing the ring size in polycatenanes leads to a higher viscosity, primarily because of the increase in polymer size. The shear thinning exponent (SI Table E.1) also increases slightly for larger rings. This is due to the additional molecular degrees of freedom, leading to improved alignment with the flow, demonstrated with decreased orientation angles in SI Fig. B.4c.
@@
At very low Wi number, short, stiff polycatenanes have decreased tumbling frequency compared to linear polymers, as shown in Fig. 7c. However, because of the interlocked rings, polycatenanes are thicker in gradient direction (as seen in SI Fig. B.3a), making them more responsive to the flow strength. As a result, a slight increase in Wi causes more tumbling in short polycatenanes compared with long polycatenanes and linear polymers, where tumbling frequency remains fairly constant in low Wi regime. For example, a polycatenane with 5 rings has increased tumbling frequency at $\mathrm{Wi}\approx 0.2$, compared to linear polymers and polycatenanes with more than 10 rings where the increase occurs around $\mathrm{Wi}\approx 1$. At moderate to high Wi number, polycatenanes tumble more frequently than linear polymers. For instance, at $\mathrm{Wi}=200$, a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$) tumbles 3 times more frequently than the linear polymer with 80 beads. Additionally, more rings in a polycatenane decreases tumbling frequency at moderate to high Wi number. This can be explained by the resulting increase in aspect ratio of the polymer, making the polymer thickness less effective, causing the chains to behave more like linear polymers. Increasing the ring size has a similar effect on tumbling frequency as increasing the number of rings, as shown in Fig. 8c. The larger more flexible rings cause the polycatenanes to behave more like linear polymers with decreased tumbling frequency.
According to Fig. 7d, polycatenanes have a $\psi_{1}$ value about two orders of magnitude lower than that of linear polymers at the same molecular weight, with the number of rings having only a minimal effect. As shown in SI Table E.2, the power-law exponent ($\beta$) for the polycatenanes is slightly higher than linear polymers of comparable molecular weight, with more rings further increasing the exponent. For example, for a polycatenane with 8 beads per ring and 10 rings ($M_{w}=80$), $\beta=-1.067$ compared with a value of -1.239 corresponding to a linear polymer with 80 beads.
@@ 3 Refer to caption
Figure 9: Comparison of linear and MIP polymers of similar molecular weight in extensional (a, b) and shear (c, d) flow. Fractional extension in the flow direction $(\langle X_{bb}\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black) has 80 backbone beads ($M_{w}=80$). The polyrotaxane (blue) has 40 backbone beads including 2 capping beads and 6 rings ($M_{w}=88$). The daisy chain (red) has 4 linear segments of 8 beads each interlocked with 6 8 bead rings ($M_{w}=74$) and includes 2 capping beads. The polycatenane (green) has 10 interconnected rings of 8 beads each $M_{w}=80$.
@@
The presence of mechanical bonds gives rise to notable quantitative differences compared with linear polymers. Unlike linear polymers that coil under weak flow, in polyrotaxanes, ring-backbone repulsion causes more extended equilibrium configurations and earlier chain extension in flow. Polyrotaxanes show lower normal stress difference and weaker elasticity. Extensional viscosity is higher in polyrotaxanes especially in shorter chains with more rings due to increased rigidity and resistance to elongation. Under shear flow, polyrotaxanes extend more in all directions and have higher viscosity, due their larger size, with reduced shear thinning originating from less alignment with flow. In addition, tumbling frequency is higher and it increases with number of rings due to enhanced hydrodynamic cross-section from threaded rings.
@@
Daisy chains extend more than linear polymers especially in weak to moderate flows and have lower normal stress difference, extensional and shear viscosity, as well as reduced shear thinning. They tumble more frequently at high Wi, due to higher gradient-direction extent and aspect ratio effects with higher density of mechanical bonds enhancing this effect, causing greater deviation from linear polymers.
@@
Polycatenanes are more extended and have lower first normal stress difference than linear chains except at very low Wi. They lack the smooth transition from coiled to stretched states typical of linear polymers. Polycatenanes show significantly lower extensional and shear viscosities, as well as weaker shear thinning. More or larger rings in polycatenanes increases viscosity, primarily due to increased size. Polycatenanes tumble more frequently, with fewer or smaller rings enhancing this effect.
@@
Overall, the interlocked topology of each MIP leads to a distinct set of rheological properties, as illustrated in Fig. B.5. Here we give an example of each polymer type, all with a similar molecular weight. We compare polymer types with similar maximum contour lengths ($L$) in SI Fig. B.5. For a given molecular weight, the polycatenane exhibits the greatest relative stretch in extensional flow (Fig. B.5a), while the polyrotaxane shows the greatest stretch in shear flow (Fig. B.5c). In shear, the daisy chain and the polycatenane have fairly consistent stretch across a wide range of Wi, varying only between 0.18-0.31 and 0.28–0.34, respectively, much less than the variability observed in linear polymers and polyrotaxanes. For a given molecular weight, the polymers also have distinct viscosity profiles. In extensional flow, the linear polymer has the greatest extensional viscosity for all but the lowest Wi range, as seen in Fig. B.5b. The mechanical bonds in all MIP types lead to lower extensional viscosity, with the daisy chain behaving similar to the linear polymer and the polyrotaxane and polycatenane showing weaker scaling of their extensional viscosity than the linear polymer and daisy chain. In shear flow (Fig. B.5d), the linear polymer has higher shear viscosity than the MIPs except the polyrotaxane, which overtakes the linear polymer at a cross-over around $\mathrm{Wi}\approx 50$. This demonstrates the reduced shear thinning of the polyrotaxane. Indeed all of the MIPs’ shear viscosity scales less strongly with Wi than the linear polymer, demonstrating that the mechanical bonds suppress shear-thinning behavior. Similar to extensional viscosity, the shear viscosity of the polycatenane is the lowest of all polymer types.
In line with the varied rheological responses that we have reported, MIPs have already shown their potential in a range of applications. Slide-ring gels, a category of polyrotaxanes in which mechanically interlocked rings function as movable cross-links, could potentially be used as protective coatings in automobiles and smartphones due to their scratch-resistance and self-healing properties Noda et al. (2014). Cyclodextrin-based polyrotaxanes have shown promise in stretchable electronics and energy storage materials. Their innovative use in stretchable conductive composite and binders for anodes in lithium-ion batteries helps to overcome critical challenges in these areas Du et al. (2024). Olympic gels (materials formed entirely from interlocked rings) Raphaël et al. (1997), recently synthesized from over 16,000 DNA rings, exhibit unique nonlinear elasticity and swelling behavior with potential applications in synthetic biology and bioengineering, including artificial micro-reactors for targeted delivery and artificial nuclei in synthetic cells Speed et al. (2024). Moreover, mechanically interlocked [c2]daisy chain structures embedded in polymer networks enable advanced shape-memory materials for use as actuators in releasing and lifting objects, with further potential in soft robots via programmed stimuli-responsive motions Zhou et al. (2024).
@@
Our results hint at additional potential uses of MIPs, as well. High molecular weight polymers are known to reduce drag in pipe flow Larson (2003); Perkins et al. (1997); Dimitropoulos et al. (1998); Paterson and Abernathy (1970), at least partially due to the suppression of vortex formation caused by extended polymer configurations. The lower shear viscosity for MIPs compared to typical linear polymers and the greater extension of polyrotaxanes, in particular, suggest that MIPs could be developed into drag reducing agents. Similarly, extended polymers in liquid petroleum-based fuels favor the formation of larger droplets that are less prone to ignition during accidental fuel releases Chao et al. (1984); Wei et al. (2015); Lhota et al. (2024). The greater extension of MIPs also suggests their potential development as mist-suppressing fuel additives to improve liquid fuel safety.
@@
As a first step toward characterizing the flow behavior of mechanically interlocked polymers, this work provides valuable insights into MIP dynamics in dilute solutions under weak to moderate flow conditions. However, strong flow regimes remain unexplored as our current model fails to maintain polymer integrity at high Wi. Under such conditions, large hydrodynamic forces can cause the rings in all three MIP architectures to cross the polymer backbone, or each other in the case of polycatenanes. Although we could change the model parameters to prevent bond crossing at very high Wi, we chose to use the standard set of FENE and WCA parameters to ensure that our results remain directly comparable with previous studies in the low to moderate Wi regime. Despite the model deficiencies at high Wi, the observed degradation is nevertheless consistent with experimental findings, which show that rotaxanes degrade via dethreading and unstoppering under strong mechanical forces Muramatsu et al. (2021); Zhang and De Bo (2019). Other experimental observations suggest that mechanical bonds in catenanes are comparable in strength to covalent bonds Lee et al. (2016). To fully capture degradation mechanisms in strong flow, future work will move beyond the FENE bond model and incorporate dissociable bonds, revealing combined chemical and mechanical modes of polymer failure. This future work will lead to a better understanding of the durability of MIPs in flow for potential engineering applications.